OpenStudy (anonymous):

Find the critical numbers of f(t) = sin^2 t - cos t

OpenStudy (cruffo):

What steps are involved in finding critical numbers?

OpenStudy (cruffo):

1.) find the first derivative 2.) find numbers that make 1st derivative = 0 or undefined did you find the 1st derivative?

OpenStudy (cruffo):

ok.. so \[f'(t) = 2\sin(t)\cos(t) + \sin(t)\]

OpenStudy (anonymous):

Sorry,internet disconnected. Let me show you where I've gotten to.

OpenStudy (anonymous):

f prime of t = cos^2 t + sin t = 1 + sin t(1 - sin t)

OpenStudy (cruffo):

You need chain rule for that first term \[\sin^2(t) = [\sin(t)]^2\]

OpenStudy (anonymous):

Oh my God! Thanks a bunch cruffo! I didn't properly differentiate sin^2(t), did I? :P

OpenStudy (cruffo):

np : )

OpenStudy (anonymous):

\[\left\{\{t\to 0\},\left\{t\to -\frac{2 \pi }{3}\right\},\left\{t\to \frac{2 \pi }{3}\right\}\right\} \]

OpenStudy (cruffo):

\[\sin(t)(2\cos(t) + 1) = 0\] sin(t) = 0 when t = any multiple of pi 2cos(t) + 1 = 0 2cos(t) = -1 cos(t) = -1/2 t = 2pi/3 + 2kpi or 4pi/3 +2kpi

OpenStudy (cruffo):

\[t = \left\{k\pi, \; \pm\frac{2\pi}{3} + 2k\pi\right\}\]